# Microlensing optical depth, event rate, and limits on compact objects in dark matter based on 20 years of OGLE observations of the Small Magellanic Cloud

### v1

@@ 4 Detected Events
A formally statistically significant signal of the microlensing parallax was also identified for OGLE-SMC-07. The fit improved by $\Delta\chi^{2}=23.5$ when including the parallax. This improvement mainly originates from the data points collected prior to the main event, in the time range $4112,458,284<\mathrm{HJD}<2,458,$ 411, during which the event was magnified by $\approx 0.05$ mag compared to the mean baseline brightness. The parallax is also marginally ($\Delta\chi^{2}=8.6$) detected in the light curve of OGLE-SMC-03, as was originally noticed by Wyrzykowski et al. (2011b). We identified four degenerate solutions for both events–two differing by the sign of $u_{0}$ (Smith et al., 2003) and two resulting from the jerk-parallax degeneracy (Gould, 2004). The parameters of both events (short timescales and large microlensing parallaxes) indicate that the lenses are located in the foreground Milky Way disk.
@@ 4.1 Comparison with Previous Searches for Microlensing Events toward the SMC
The analyzed data set, by construction, contains the majority of OGLE-III light curves that were previously studied by Wyrzykowski et al. (2011b), who reported the discovery of three microlensing events. We were able to successfully recover two of these events (OGLE-SMC-03 and OGLE-SMC-04). Their best-fit parameters are very similar to those measured by Wyrzykowski et al. (2011b) (Table 5). However, one event (OGLE-SMC-02) was not selected by our automated pipeline because it did not pass all selection cuts, in particular those related to the PSPL fit quality. That is not surprising because this event is known to be a weak (non-caustic-crossing) binary-lens event whose light curve significantly deviates from a simple PSPL model (Dong et al., 2007). For completeness, we include the light curve of OGLE-SMC-02 in the upper left panel of Figure 2.
@@ 5 Number of Source Stars
We found that these correction factors were very similar to those derived in Paper 1 for the LMC. This conclusion is further supported by analyzes of simulations carried out by Wyrzykowski et al. (2011a) and Wyrzykowski et al. (2011b), which show that the correction factors for the LMC and SMC are consistent within 5%. Because the number of SMC fields included in the Holtzman et al. (2006) database is much smaller than the number of LMC fields (64), we adopted the correction factors derived in Paper 1.
We then estimated the surface density of foreground stars by fitting the following formula
@@
Figure 5: Example detection efficiency curves for representative fields SMC726 and SMC734. The solid, dashed, and dash-dotted lines represent detection efficiencies averaged for sources brighter than $I=21$, $I=21.5$, and $I=22$ mag, respectively.
@@ 6 Detection Efficiency
The parameters of events were drawn from physically motivated distributions. The time of the maximum was sampled from a uniform distribution within the time ranges $0002,452,000\leq t_{0}\leq 2,459,$ 000 (for objects observed during both OGLE-III and OGLE-IV) and $0002,455,000\leq t_{0}\leq 2,459,$ 000 (for stars observed during OGLE-IV only). The impact parameters were drawn from a uniform distribution from the range $0\leq u_{0}\leq 1$, while the Einstein timescales were drawn from a log-uniform distribution from the range $1\leq t_{\rm E}\leq 10^{4}$ days. For the blending parameter $f_{\rm s}$, we used an empirical distribution created by matching stars detected on OGLE reference images with those detected by HST.
@@ 7 Optical Depth and Event Rate
For our calculations, we use the five events that met all selection criteria outlined in Section 3, alongside the number of sources brighter than $I=22$ mag ($N_{\rm s}=(12.4\pm 1.4)\times 10^{6}$) and the detection efficiencies calculated for sources brighter than this limit. The total microlensing optical depth toward the OGLE-IV SMC fields is found to be $\tau=(0.32\pm 0.18)\times 10^{-7}$, while the event rate is determined to be $\Gamma=(1.18\pm 0.57)\times 10^{-7}\,\mathrm{yr}^{-1}\,\mathrm{star}^{-1}$. The uncertainties of these quantities were calculated using methods proposed by Han & Gould (1995) and Mróz et al. (2019a). These error bars reflect only the uncertainties caused by the small number of events and by the uncertainties in Einstein timescales. There is an additional source of error (approximately 11%) related to the accuracy of measuring the number of source stars.
@@ 8 Expected Number of Events from Stellar Populations
which was first introduced by Batista et al. (2011). Here, $M$ is the mass of the lens, $f(\boldsymbol{\mu})$ is the two-dimensional probability function for a given source–lens relative proper motion $\boldsymbol{\mu}$, $\theta_{\rm E}$ is the angular Einstein radius, $\nu(D_{\rm l})$ is the number density of lenses at a given distance $D_{\rm l}$, and $g(M)$ is the mass function of lenses. The total event rate in a given direction can be expressed as the following integral
[…]
where $\rho(D_{\rm s})$ is the density of sources, and the coefficient $\beta$ quantifies how the number of observed sources changes with their distance (Kiraga & Paczyński, 1994). It is generally assumed that $0\leq\beta\leq 1$. Here we adopt $\beta=0$ as a fiducial value, but we demonstrate below that the results weakly depend on the choice of this parameter.
The integrals in Equation (7) are evaluated numerically using the Monte Carlo integration technique. For example, to calculate the multidimensional integral in the numerator of Equation (7), we draw a random sample of $10^{8}$ points from a 5-dimensional cube $D_{\rm l,min}\leq D_{\rm l}\leq D_{\rm l,max}$, $D_{\rm s,min}\leq D_{\rm s}\leq D_{\rm s,max}$, $M_{\rm min}\leq M\leq M_{\rm max}$, $\mu_{\rm min}\leq\mu_{x,y}\leq\mu_{\rm max}$, evaluate the integrand at each point, and then calculate its mean value. To speed up the calculations, we change from the variable $M$ to $\log M$. All integrals are evaluated with precision better than 1%.
As a by-product of the Monte Carlo integration, we also evaluate the expected event timescale distribution $d\Gamma/d\log t_{\rm E}$ (where $t_{\rm E}=\theta_{\rm E}/\mu$). The total number of expected events in a given field can be found by
@@ 8.1 SMC Self-Lensing Models
We analyze two SMC self-lensing models. The first one is based on the model proposed by Calchi Novati et al. (2013), in which the total stellar mass (within 5 kpc of the SMC center) is taken to be $M_{*}=1.0\times 10^{9}\,M_{\odot}$ (Bekki & Chiba, 2009). However, this quantity is by no means well determined, and Calchi Novati et al. (2013) argued that $M_{*}$ is known with an accuracy of up to a factor of 2. For example, McConnachie (2012) reports $M_{*}=4.6\times 10^{8}\,M_{\odot}$. Because the derived microlensing event rate is directly proportional to the total stellar mass, the former quantity is also accurate up to a factor of 2.
@@
For both stellar distribution models, we use the empirical model of the SMC kinematics that was derived using the Gaia DR3 data (Appendix A). We assume that the mass function of lenses is identical to that measured by Kroupa (2001). For the evaluation of integrals in Equation (7), we adopt $D_{\rm l,min}=D_{\rm s,min}=48$ kpc, $D_{\rm l,max}=D_{\rm s,max}=76$ kpc, $\mu_{\rm min}=-0.5$ mas yr-1, $\mu_{\rm max}=0.5$ mas yr-1, $M_{\rm min}=0.01\,M_{\odot}$, and $M_{\rm max}=10\,M_{\odot}$. The event rate was evaluated at the center of each field (Table 2). For four central fields (SMC719, SMC720, SMC725, and SMC726), the integrals in Equation (7) were evaluated at the centers of all 32 subfields (that is, separately for each CCD detector), and then averaged using the expected number of sources (denominator in Equation (7)) as a weight.
@@ 8.3 Comparison with Observations
where the summation is performed over all observed fields, $N_{{\rm s},i}$ is the number of source stars brighter than $I=22$ mag in a given field, and $\Delta T_{i}$ is the duration of its observations. Self-lensing events are expected to have relatively long timescales, with the average Einstein timescale is $\langle t_{\rm E}\rangle\approx 110$ days, primarily because of low relative lens–source proper motions. In contrast, events with Milky Way lenses tend to have much shorter timescales due to higher relative proper motions, leading to an average of $\langle t_{\rm E}\rangle\approx 30$ days.
The comparison between the observed cumulative distribution of event timescales and the predictions from our models is presented in Figure 8. The expected distributions match the observations reasonably well. To quantify that, we performed a Kolmogorov-Smirnov test, resulting in $p$-values of 0.957 and 0.965 for the models including Han & Gould (2003) and Cautun et al. (2020) Milky Way disk models, respectively (for $\beta=0$ and the Calchi Novati et al. (2013) SMC model). We found similar $p$-values for models with $\beta=1$ and those calculated using the Mróz & Poleski (2018) approach. The results of these tests confirm that there are not any statistically significant deviations between the observed and expected timescales.
Figure 7: Expected distributions of observed event timescales of microlensing events from known stellar populations in the SMC (Calchi Novati et al. 2013; blue line) and the foreground Milky Way disk (red line). Left panel—assuming the Milky Way model of Han & Gould (2003), right panel—Cautun et al. (2020).
@@ 9 Limits on Primordial Black Holes in Dark Matter
Figure 9 shows the total expected number of microlensing events, assuming that the entire Milky Way dark matter halo consists of compact objects of identical mass $M$. We expect the largest number of events (approximately 233) for $M=0.013\,M_{\odot}$. We should have detected about 103 events if entire dark matter was made of $1\,M_{\odot}$ PBHs, 47 events—$10\,M_{\odot}$, and 18 events—$1000\,M_{\odot}$. That should be compared to five events in our statistical sample (Section 3).
@@
We calculated the 95% upper limits on $f$ using the same methodology as in Paper 2. We present two types of limits. For the strict limits, we assumed that all five microlensing events in our statistical sample are due to known stellar populations in the SMC or in the Milky Way disk. These limits are marked with a solid red line in Figure 10. When we relax this assumption and allow a possibility of a PBH origin for these microlensing events, the inferred limits become slightly weaker. They are shown by red dashed and dotted lines in Figure 10. The relaxed limits slightly depend on the choice of the Milky Way disk or the SMC model. In Figure 10 we present two example results for the Han & Gould (2003) (dashed line) and Cautun et al. (2020) (dotted line) Milky Way disk models.
@@ 10.2 Limits on Compact Objects as Dark Matter
The inferred limits depend on the choice of the Milky Way halo model. In this study, we used the contracted halo model from Cautun et al. (2020), which was fitted to the Milky Way’s rotation curve derived from the Gaia DR2 data (Eilers et al., 2019). However, several recent works (e.g., Wang et al., 2023; Jiao et al., 2023; Ou et al., 2024) have indicated evidence for a declining Milky Way rotation curve at Galactocentric distances greater than 15–20 kpc, based on the more recent Gaia DR3 data. This suggests that the estimated mass of the Milky Way halo may be smaller than previously thought. These new findings, however, are in tension with other mass estimates derived from observations of globular clusters, dwarf satellites, and streams. Koop et al. (2024) pointed out that—at large Galactocentric radii—disequilibrium and deviations from axisymmetry could lead to large systematic errors in the inferred rotation curve.
@@
Additionally, Green (2025) investigated how the choice of the halo model impacts microlensing constraints on the abundance of PBHs. They considered two power law halo models (Evans, 1994) from Alcock et al. (1996) (models B and C) and found that the tightest limit on $f$ is a factor of 1.7 (1.5) lower (higher) for model B (C) compared to the standard halo model. We note, however, that none of these models accurately fit the Milky Way rotation curve (Mróz et al., 2019b): model B consistently predicts higher rotation speeds, while model C predicts lower speeds than actually observed. In contrast, the rotation curves predicted by the models from Cautun et al. (2020), Jiao et al. (2023), and Ou et al. (2024) match the data very well. We conclude that the uncertainties on $f$ associated with the choice of the Milky Way halo model are substantially smaller than those indicated by the predictions of the B and C halo models of Alcock et al. (1996).
Recently Li et al. (2025) studied the constraints on the abundance of planetary-mass PBHs based on a sample of 2617 events detected by OGLE in the direction of the Galactic bulge (Mróz et al., 2017). They investigated the dependence of these constraints on the choice of dark matter halo models and found that the constraints derived assuming the Einasto and Burkert profiles are much weaker than those from the standard Navarro-Frenk-White profile. We check the same models to study the robustness of our results.
The Einasto profile used by Li et al. (2025) was taken from Hertzberg et al. (2021), who in turn used the dark matter profile from Pieri et al. (2011). The latter model was found from the Milky Way Aquarius simulations by Springel et al. (2008). Unfortunately, the parameters of the model changed during this telephone game: Pieri et al. (2011) reported the normalization of the profile of $\rho_{s}=2.8\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-1}$, while Li et al. (2025) used $\rho_{s}=2.08\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-1}$ (thus, the constrains published by Li et al. (2025) are underestimated by 35%). Assuming that the parameters from Pieri et al. (2011) are correct, the microlensing optical depth toward the SMC center is $\tau=11.1\times 10^{-7}$, i.e., a factor 1.7 larger than in our fiducial model. However, the Milky Way rotation curve predicted by the former model does not match the data (Mróz et al., 2019b). We also found that the rotation curves calculated using Burkert profiles considered in Li et al. (2025) do not match the observed one at all (Mróz et al., 2019b). That is not surprising as these models come from Donato et al. (2009) and they were intended to fit the weak gravitational lensing observations of some galaxies, but not the Milky Way dark matter profile.

### v2

@@ 4 Detected Events
A formally statistically significant signal of the microlensing parallax was also identified for OGLE-SMC-07. The fit improved by $\Delta\chi^{2}=23.5$ when including the parallax. This improvement mainly originates from the data points collected prior to the main event, in the time range $2,458,284<\mathrm{HJD}<2,458,411$, during which the event was magnified by $\approx 0.05$ mag compared to the mean baseline brightness. The parallax is also marginally ($\Delta\chi^{2}=8.6$) detected in the light curve of OGLE-SMC-03, as was originally noticed by Wyrzykowski et al. (2011b). We identified four degenerate solutions for both events—two differing by the sign of $u_{0}$ (Smith et al., 2003) and two resulting from the jerk-parallax degeneracy (Gould, 2004). The parameters of both events (short timescales and large microlensing parallaxes) indicate that the lenses are located in the foreground Milky Way disk.
@@ 4.1 Comparison with Previous Searches for Microlensing Events toward the SMC
The analyzed data set, by construction, contains the majority of OGLE-III light curves that were previously studied by Wyrzykowski et al. (2011b), who reported the discovery of three microlensing events. We were able to successfully recover two of these events (OGLE-SMC-03 and OGLE-SMC-04). Their best-fit parameters are very similar to those measured by Wyrzykowski et al. (2011b) (Table 5). However, one event (OGLE-SMC-02) was not selected by our automated pipeline because it did not pass all selection cuts, in particular those related to the PSPL fit quality. That is not surprising because this event is known to be a weak (non-caustic-crossing) binary-lens event whose light curve significantly deviates from a simple PSPL model (Dong et al., 2007). For completeness, we include the light curve of OGLE-SMC-02 in the upper-left panel of Figure 2.
@@ 5 Number of Source Stars
We found that these correction factors were very similar to those derived in Paper 1 for the LMC. This conclusion is further supported by analyses of simulations carried out by Wyrzykowski et al. (2011a) and Wyrzykowski et al. (2011b), which show that the correction factors for the LMC and SMC are consistent within 5%. Because the number of SMC fields included in the Holtzman et al. (2006) database is much smaller than the number of LMC fields (64), we adopted the correction factors derived in Paper 1.
We then estimated the surface density of foreground stars by fitting the following formula:
@@
Figure 5: Example detection efficiency curves for representative fields SMC726 and SMC734. The solid, dashed, and dash–dotted lines represent detection efficiencies averaged for sources brighter than $I=21$, $I=21.5$, and $I=22$ mag, respectively.
@@ 6 Detection Efficiency
The parameters of events were drawn from physically motivated distributions. The time of the maximum was sampled from a uniform distribution within the time ranges $2,452,000\leq t_{0}\leq 2,459,000$ (for objects observed during both OGLE-III and OGLE-IV) and $2,455,000\leq t_{0}\leq 2,459,000$ (for stars observed during OGLE-IV only). The impact parameters were drawn from a uniform distribution from the range $0\leq u_{0}\leq 1$, while the Einstein timescales were drawn from a log-uniform distribution from the range $1\leq t_{\rm E}\leq 10^{4}$ days. For the blending parameter $f_{\rm s}$, we used an empirical distribution created by matching stars detected on OGLE reference images with those detected by HST.
@@ 7 Optical Depth and Event Rate
For our calculations, we use the five events that met all selection criteria outlined in Section 3, alongside the number of sources brighter than $I=22$ mag ($N_{\rm s}=(12.4\pm 1.4)\times 10^{6}$) and the detection efficiencies calculated for sources brighter than this limit. The total microlensing optical depth toward the OGLE-IV SMC fields is found to be $\tau=(0.32\pm 0.18)\times 10^{-7}$, while the event rate is determined to be $\Gamma=(1.18\pm 0.57)\times 10^{-7}\,\mathrm{yr}^{-1}\,\mathrm{star}^{-1}$. The uncertainties of these quantities were calculated using methods proposed by Han & Gould (1995) and Mróz et al. (2019a). These error bars reflect only the uncertainties caused by the small number of events and by the uncertainties in Einstein timescales. There is an additional source of error (approximately 11%) related to the accuracy of measuring the number of source stars. The contribution of each event to the total optical depth and event rate is detailed in Table 7.
Table 7: Microlensing Optical Depth and Event Rate Toward the LMC
Event $\langle\varepsilon(t_{\rm E})\rangle$ $\langle 1/\varepsilon(t_{\rm E})\rangle$ $\langle t_{\rm E}/\varepsilon(t_{\rm E})\rangle$ $\tau$ $\Gamma$
(days) ($10^{-7}$) $(10^{-7}\,\mathrm{yr}^{-1})$
OGLE-SMC-03 0.1848 5.430 235.76 0.0428 0.2294
OGLE-SMC-04 0.1205 8.379 156.17 0.0284 0.3540
OGLE-SMC-05 0.2635 3.827 439.51 0.0799 0.1617
OGLE-SMC-06 0.3599 2.783 787.51 0.1431 0.1176
OGLE-SMC-07 0.1336 7.593 153.36 0.0279 0.3208
Total: $0.32\pm 0.18$ $1.18\pm 0.57$
@@ 8 Expected Number of Events from Stellar Populations
which was first introduced by Batista et al. (2011). Here, $M$ is the mass of the lens, $f(\boldsymbol{\mu})$ is the 2D probability function for a given source–lens relative proper motion $\boldsymbol{\mu}$, $\theta_{\rm E}$ is the angular Einstein radius, $\nu(D_{\rm l})$ is the number density of lenses at a given distance $D_{\rm l}$, and $g(M)$ is the mass function of lenses. The total event rate in a given direction can be expressed as the following integral:
$\Gamma=\frac{\int_{0}^{\infty}\rho(D_{\rm s})D_{\rm s}^{2-2\beta}dD_{\rm s}\int_{0}^{D_{\rm s}}\frac{d^{4}\Gamma}{dD_{\rm l}dMd^{2}\mu}dD_{\rm l}dMd^{2}\mu}{\int_{0}^{\infty}\rho(D_{\rm s})D_{\rm s}^{2-2\beta}dD_{\rm s}},$ (7)
where $\rho(D_{\rm s})$ is the density of sources, and the coefficient $\beta$ quantifies how the number of observed sources changes with their distance (Kiraga & Paczyński, 1994). It is generally assumed that $0\leq\beta\leq 1$. Here, we adopt $\beta=0$ as a fiducial value, but we demonstrate below that the results weakly depend on the choice of this parameter.
The integrals in Equation (7) are evaluated numerically using the Monte Carlo integration technique. For example, to calculate the multidimensional integral in the numerator of Equation (7), we draw a random sample of $10^{8}$ points from a 5D cube $D_{\rm l,min}\leq D_{\rm l}\leq D_{\rm l,max}$, $D_{\rm s,min}\leq D_{\rm s}\leq D_{\rm s,max}$, $M_{\rm min}\leq M\leq M_{\rm max}$, $\mu_{\rm min}\leq\mu_{x,y}\leq\mu_{\rm max}$, evaluate the integrand at each point, and then calculate its mean value. To speed up the calculations, we change from the variable $M$ to $\log M$. All integrals are evaluated with precision better than 1%.
As a byproduct of the Monte Carlo integration, we also evaluate the expected event timescale distribution $d\Gamma/d\log t_{\rm E}$ (where $t_{\rm E}=\theta_{\rm E}/\mu$). The total number of expected events in a given field can be found by
@@ 8.1 SMC Self-Lensing Models
We analyze two SMC self-lensing models. The first one is based on the model proposed by Calchi Novati et al. (2013), in which the total stellar mass (within 5 kpc of the SMC center) is taken to be $M_{*}=1.0\times 10^{9}\,M_{\odot}$ (Bekki & Chiba, 2009). However, this quantity is by no means well determined, and Calchi Novati et al. (2013) argued that $M_{*}$ is known with an accuracy of up to a factor of 2. For example, McConnachie (2012) reported $M_{*}=4.6\times 10^{8}\,M_{\odot}$. Because the derived microlensing event rate is directly proportional to the total stellar mass, the former quantity is also accurate up to a factor of 2.
@@
For both stellar distribution models, we use the empirical model of the SMC kinematics that was derived using the Gaia DR3 data (see the Appendix A). We assume that the mass function of lenses is identical to that measured by Kroupa (2001). For the evaluation of integrals in Equation (7), we adopt $D_{\rm l,min}=D_{\rm s,min}=48$ kpc, $D_{\rm l,max}=D_{\rm s,max}=76$ kpc, $\mu_{\rm min}=-0.5$ mas yr-1, $\mu_{\rm max}=0.5$ mas yr-1, $M_{\rm min}=0.01\,M_{\odot}$, and $M_{\rm max}=10\,M_{\odot}$. The event rate was evaluated at the center of each field (Table 2). For four central fields (SMC719, SMC720, SMC725, and SMC726), the integrals in Equation (7) were evaluated at the centers of all 32 subfields (that is, separately for each CCD detector), and then averaged using the expected number of sources (denominator in Equation (7)) as a weight.
@@ 8.3 Comparison with Observations
where the summation is performed over all observed fields, $N_{{\rm s},i}$ is the number of source stars brighter than $I=22$ mag in a given field, and $\Delta T_{i}$ is the duration of its observations. Self-lensing events are expected to have relatively long timescales, with the average Einstein timescale being $\langle t_{\rm E}\rangle\approx 110$ days, primarily because of low relative lens–source proper motions. In contrast, events with Milky Way lenses tend to have much shorter timescales due to higher relative proper motions, leading to an average of $\langle t_{\rm E}\rangle\approx 30$ days.
The comparison between the observed cumulative distribution of event timescales and the predictions from our models is presented in Figure 8. The expected distributions match the observations reasonably well. To quantify that, we performed a Kolmogorov–Smirnov test, resulting in $p$-values of 0.957 and 0.965 for the models including Han & Gould (2003) and Cautun et al. (2020) Milky Way disk models, respectively (for $\beta=0$ and the Calchi Novati et al. (2013) SMC model). We found similar $p$-values for models with $\beta=1$ and those calculated using the Mróz & Poleski (2018) approach. The results of these tests confirm that there are not any statistically significant deviations between the observed and expected timescales.
Figure 7: Expected distributions of observed event timescales of microlensing events from known stellar populations in the SMC (Calchi Novati et al. 2013; blue line) and the foreground Milky Way disk (red line). Left panel: assuming the Milky Way model of Han & Gould (2003); right panel: Cautun et al. (2020).
@@ 9 Limits on Primordial Black Holes in Dark Matter
[…]
@@
We calculated the 95% upper limits on $f$ using the same methodology as in Paper 2. We present two types of limits. For the strict limits, we assumed that all five microlensing events in our statistical sample are due to known stellar populations in the SMC or in the Milky Way disk. These limits are marked with a solid red line in Figure 10. When we relax this assumption and allow for a possibility of a PBH origin for these microlensing events, the inferred limits become slightly weaker. They are shown by red dashed and dotted lines in Figure 10. The relaxed limits slightly depend on the choice of the Milky Way disk or the SMC model. In Figure 10 we present two example results for the Han & Gould (2003) (dashed line) and Cautun et al. (2020) (dotted line) Milky Way disk models.
@@ 10.2 Limits on Compact Objects as Dark Matter
The inferred limits depend on the choice of the Milky Way halo model. In this study, we used the contracted halo model from Cautun et al. (2020), which was fitted to the Milky Way’s rotation curve derived from the Gaia DR2 data (Eilers et al., 2019). However, several recent works (e.g., Wang et al., 2023; Jiao et al., 2023; Ou et al., 2024) have indicated evidence for a declining Milky Way rotation curve at Galactocentric distances greater than 15–20 kpc, based on the more recent Gaia DR3 data. This suggests that the estimated mass of the Milky Way halo may be smaller than previously thought. These new findings, however, are in tension with other mass estimates derived from observations of globular clusters, dwarf satellites, and streams. Koop et al. (2024) pointed out that—at large Galactocentric radii—disequilibrium and deviations from axisymmetry could lead to large systematic errors in the inferred rotation curves.
@@
Additionally, Green (2025) investigated how the choice of the halo model impacts microlensing constraints on the abundance of PBHs. They considered two power-law halo models (Evans, 1994) from Alcock et al. (1996) (models B and C) and found that the tightest limit on $f$ is a factor of 1.7 (1.5) lower (higher) for model B (C) compared to the standard halo model. We note, however, that none of these models accurately fit the Milky Way rotation curve (Mróz et al., 2019b): model B consistently predicts higher rotation speeds, while model C predicts lower speeds than actually observed. In contrast, the rotation curves predicted by the models from Cautun et al. (2020), Jiao et al. (2023), and Ou et al. (2024) match the data very well. We conclude that the uncertainties on $f$ associated with the choice of the Milky Way halo model are substantially smaller than those indicated by the predictions of the B and C halo models of Alcock et al. (1996).
Recently, Li et al. (2025) studied the constraints on the abundance of planetary-mass PBHs based on a sample of 2617 events detected by OGLE in the direction of the Galactic bulge (Mróz et al., 2017). They investigated the dependence of these constraints on the choice of dark matter halo models and found that the constraints derived assuming the Einasto and Burkert profiles are much weaker than those from the standard Navarro–Frenk–White profile. We check the same models to study the robustness of our results.
The Einasto profile used by Li et al. (2025) was taken from Hertzberg et al. (2021), who in turn used the dark matter profile from Pieri et al. (2011). The latter model was found from the Milky Way Aquarius simulations by Springel et al. (2008). Unfortunately, the parameters of the model changed during this telephone game: Pieri et al. (2011) reported the normalization of the profile of $\rho_{s}=2.8\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-1}$, while Li et al. (2025) used $\rho_{s}=2.08\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-1}$ (thus, the constrains published by Li et al. (2025) are underestimated by 35%). Assuming that the parameters from Pieri et al. (2011) are correct, the microlensing optical depth toward the SMC center is $\tau=11.1\times 10^{-7}$, i.e., a factor 1.7 larger than in our fiducial model. However, the Milky Way rotation curve predicted by the former model does not match the data (Mróz et al., 2019b). We also found that the rotation curves calculated using Burkert profiles considered in Li et al. (2025) do not match the observed one at all (Mróz et al., 2019b). That is not surprising, as these models come from Donato et al. (2009) and they were intended to fit the weak gravitational lensing observations of some galaxies, but not the Milky Way dark matter profile.